if i get 30% off of $68dollars how much do i pay
Found 2 solutions by jim_thompson5910, General_Lee87:
Answer by jim_thompson5910(35256)

You save: 0.3*68 = 20.40 dollars

So in the end, your total bill is 68 - 20.40 = 47.60 dollars

-------------------------------------------------------
Alternatively,

if you get 30% off, you're really only paying for 100% - 30% = 70% of the actual price, which is 0.7*68 = 47.60 dollars

Answer by General_Lee87(58)

Origional Question: If I get 30% off of $68 how much do i pay?

Solution: 68 x .30 = 20.40
68 - 20.40 = 47.6

Answer: $47.60


First you must mulitiply 60 dollars by 30 percent to get the total amount that you recieve off. 30 percent turns into .30 because you must move the decimal point over two spaces to turn it into a workable number. Next you subtract the amount off, $20.40, from the origional amount, $60, to get the price you pay, $47.60. I hope this helped you. Best wishes and good luck.